# A Review of Design Concerns in Superconducting Quantum Circuits

### v4

@@ V.5.3 Black-box quantization
Lastly, the BBQ approach models the Josephson junction as a small nonlinear perturbation within a network of linear inductors and capacitors, treating the electromagnetic environment as a “black box” characterized by its impedance [123]. The zeros and poles of the impedance are analyzed to derive Hamiltonian parameters, such as mode frequencies and couplings. This method can also handle multimode systems involving multiple nonlinear modes, making it suitable for complex circuits where the environment significantly affects qubit dynamics. While BBQ provides a comprehensive tool for extracting key Hamiltonian parameters, it requires both the numerical impedance and the junction field distribution. These typically cannot be extracted from the same FEM solver, meaning BBQ requires at least 2 simulation steps. This makes it it computationally demanding and dependent on accurate impedance calculations.
@@ V.5.4 Example of simulation analysis
The choice of simulation analysis method depends on the specific requirements of the design and the level of precision needed. Each method has its advantages and limitations, and sometimes a combination of methods is employed to leverage their respective strengths. To demonstrate how to perform quantum analysis of FEM simulations, we take the example of the transmon-cavity system described earlier (Section II). We use Ansys as the FEM solver and employ the sub-section simulation methodology. This approach allows us to simulate the transmon and resonator separately, greatly speeding the simulations and making it easy to customize either structure separately. We include the resonator coupling capacitor in both simulations, overlapping the sections to ensure that no inter-section couplings are missed. For the transmon section, we run an electrostatic simulation using Ansys Q3D to obtain the transmon’s capacitance matrix, including the coupling capacitor $C_{C}$ connecting the qubit and the resonator. Using the capacitance matrix and a design value for the junction $E_{J}$, we can then calculate the transmon frequency and anharmonicity. Similarly, we simulate the resonator subsystem using the Ansys HFSS solver to run an eigenmodal simulation. This gives both the resonant frequency $\omega_{r}$ and linewidth $\kappa$, with the latter determined by coupling to a feedline that we terminate with lumped-element $50\ \Omega$ resistors in the simulation. The effective resonator capacitance $C_{r}$ is calculated using $C_{r}=\frac{\pi}{m\omega_{r}Z_{c}}$, where $Z_{c}$ is the characteristic impedance of the waveguide and $m$ is 2 or 4 for half-wave or quarter-wave designs, respectively. The effective inductance is then $L_{r}=(C_{r}\omega_{r}^{2})^{-1}$. Using this lumped-element model, we can then calculate transmon-resonator coupling $g$ either numerically or approximately using the Section II. Similarly, we compute the self-Kerr and cross-Kerr terms either numerically or using Eq. 10 by combining the results from the EM simulations and the quantum parameters computed from them. Note that we could have used EPR analysis to analyze the whole system and immediately extract all energies (including Kerr terms). This would have required an Ansys HFSS eigenmodal simulation of the entire transmon-resonator system, which in our experience can be done accurately but takes much longer.
@@ VI From Design to Fabrication
The updated design is re-simulated to verify improvements and ensure that no new issues have been introduced. This iterative process (shown in Fig 11) continues until the design satisfies all specifications and is considered “fab-ready.” Before finalizing for fabrication, the design undergoes a Design Rule Check (DRC)to ensure it complies with fabrication constraints such as minimum feature sizes, spacing, layer overlap, etc. Any errors detected during the DRC stage must be addressed to prevent fabrication failures. After fabrication, the device is packaged, cooled, and measured. Hamiltonian parameters are extracted to validate the simulation pipeline. Discrepancies between the measured and simulated results can arise due to inaccuracies in material models, parasitic effects, or variations during the fabrication process. These discrepancies are analyzed by comparing the required and observed Hamiltonian parameters, leading to refinements in the simulation models by adjusting material properties, re-calibrating simulation hyperparameters, or revising geometric tolerances. This feedback loop is crucial to developing robust simulation workflows and ensuring that fabricated devices perform as expected.

### v5

@@ V.5.3 Black-box quantization
Lastly, the BBQ approach models the Josephson junction as a small nonlinear perturbation within a network of linear inductors and capacitors, treating the electromagnetic environment as a “black box” characterized by its impedance [139]. The zeros and poles of the impedance are analyzed to derive Hamiltonian parameters, such as mode frequencies and couplings. This method can also handle multimode systems involving multiple nonlinear modes, making it suitable for complex circuits where the environment significantly affects qubit dynamics. While BBQ provides a comprehensive tool for extracting key Hamiltonian parameters, it requires both the numerical impedance and the junction field distribution. These typically cannot be extracted from the same FEM solver, meaning BBQ requires at least two coordinated simulation steps. This makes it it computationally demanding and dependent on accurate impedance calculations.
@@ V.5.4 Example of simulation analysis
The choice of simulation analysis method depends on the specific requirements of the design and the level of precision needed. Each method has its advantages and limitations, and sometimes a combination of methods is employed to leverage their respective strengths. To demonstrate how to perform quantum analysis of FEM simulations, we take the example of the transmon-cavity system described earlier (Section II). We use Ansys as the FEM solver and employ the sub-section simulation methodology. This approach allows us to simulate the transmon and resonator separately, greatly speeding the simulations and making it easy to customize either structure separately. We include the resonator coupling capacitor in both simulations, overlapping the sections to ensure that no inter-section couplings are missed. For the transmon section, we run an electrostatic simulation using Ansys Q3D to obtain the transmon’s capacitance matrix, including the coupling capacitor $C_{C}$ connecting the qubit and the resonator. Using the capacitance matrix and a design value for the junction $E_{J}$, we can then calculate the transmon frequency and anharmonicity. Similarly, we simulate the resonator subsystem using the Ansys HFSS solver to run an eigenmode simulation. This gives both the resonant frequency $\omega_{r}$ and linewidth $\kappa$, with the latter determined by coupling to a feedline that we terminate with lumped-element $50\ \Omega$ resistors in the simulation. The effective resonator capacitance $C_{r}$ is calculated using $C_{r}=\frac{\pi}{m\omega_{r}Z_{c}}$, where $Z_{c}$ is the characteristic impedance of the waveguide and $m$ is 2 or 4 for half-wave or quarter-wave designs, respectively. The effective inductance is then $L_{r}=(C_{r}\omega_{r}^{2})^{-1}$. Using this lumped-element model, we can then calculate transmon-resonator coupling $g$ either numerically or approximately using the Section II.3. Similarly, we compute the self-Kerr and cross-Kerr terms either numerically or using Eq. 10 by combining the results from the EM simulations and the quantum parameters computed from them. Note that we could have used EPR analysis to analyze the whole system and immediately extract all energies (including Kerr terms). This would have required an Ansys HFSS eigenmode simulation of the entire transmon-resonator system, which in our experience can be done accurately but takes much longer.
@@ VI From Design to Fabrication
The updated design is re-simulated to verify improvements and ensure that no new issues have been introduced. This iterative process (shown in Fig 11) continues until the design satisfies all specifications and is considered “fab-ready.” Before finalizing for fabrication, the design undergoes a Design Rule Check (DRC) to ensure it complies with fabrication constraints such as minimum feature sizes, spacing, layer overlap, etc. Any violations flagged during the DRC stage must be reviewed and either corrected or explicitly waived with justification, to avoid unintended fabrication failures. After fabrication, the device is packaged, cooled, and measured as described in Section VII. Hamiltonian parameters are extracted to validate the simulation pipeline. Discrepancies between the measured and simulated results can arise due to inaccuracies in material models, parasitic effects, or variations during the fabrication process. These discrepancies are analyzed by comparing the required and observed Hamiltonian parameters, leading to refinements in the simulation models by adjusting material properties, re-calibrating simulation hyperparameters, or revising geometric tolerances. This feedback loop is crucial to developing robust simulation workflows and ensuring that fabricated devices perform as expected.